# On the Dynamic Regret of Following the Regularized Leader: Optimism with History Pruning

### v1

@@ Appendix F Numerical Examples
The implementation code for the algorithms, along with the code to reproduce all experiments, is available at the following repository: (Mhaisen, 2025).
Since the FTRL and OMD variants used in our experiments neither assume prior knowledge of $P_{T}$ nor attempt to estimate it online, we compare them to OptFPRL using the tuning strategy described in Sec. $3.1$ to ensure a fair evaluation. In scenarios where no predictions are used, the predicted functions fed to the algorithms are set to zero. This allows us to understand their performance initially, independent of prediction quality. Even without predictions, the proposed algorithm outperforms the two benchmarks in many scenarios, demonstrating its performance in dynamic environments.

### v2

@@ Appendix F Numerical Examples
The implementation code for the algorithms, along with the code to reproduce all experiments, is available at the following repository: (Mhaisen, 2025). Since the FTRL and OMD variants used in our experiments neither assume prior knowledge of $P_{T}$ nor attempt to estimate it online, we compare them to OptFPRL using the tuning strategy described in Sec. $3.1$ to ensure a fair evaluation. In scenarios where no predictions are used, the predicted functions fed to the algorithms are set to zero. This allows us to understand their performance initially, independent of prediction quality. Even without predictions, the proposed algorithm outperforms the two benchmarks in many scenarios, demonstrating its performance in dynamic environments.